# Navigating the Impact of Structured Output Format on Large Language Models through the Compass of Causal Inference

### v1

@@ Abstract
Structured output from large language models (LLMs) has enhanced efficiency in processing generated information and is increasingly adopted in industrial applications. Prior studies have investigated the impact of structured output on LLMs’ generation quality, often presenting one-way findings. Some suggest that structured format enhances completeness and factual accuracy, while others argue that it restricts the reasoning capacity of LLMs and leads to reductions in standard evaluation metrics. Potential limitations of these assessments include restricted testing scenarios, weakly controlled comparative settings, and reliance on coarse metrics. In this work, we present a refined analysis using causal inference. Based on one assumed and two guaranteed constraints, we derive five potential causal structures characterizing the influence of structured output on LLMs’ generation: (1) collider without m-bias, (2) collider with m-bias, (3) single cause from instruction, (4) single cause from output format, and (5) independence. Across seven public and one developed reasoning tasks, we find that coarse metrics report positive, negative, or neutral effects of structured output on GPT-4o’s generation. However, causal inference reveals no causal impact in 43 out of 48 scenarios. In the remaining 5, 3 involve multifaceted causal structures influenced by concrete instructions.
Navigating the Impact of Structured Output Format on Large Language Models through the Compass of Causal Inference
@@ 1 Introduction
Structured output, defined as organizing generated content into a specific schema (Guo et al., 2025; OpenAI, ), has been increasingly supported by LLMs, enabling them to directly output information in user-specified formats such as JSON, XML, and YAML [^1] JSON: JavaScript Object Notation; XML: Extensible Markup Language; YAML: Yet Another Markup Language (Tang et al., 2024; Wu et al., 2024; Guo et al., 2024). This development addresses a key limitation of earlier LLMs, which required users to either fine-tune models for structured generation (Chan et al., 2024; Tang et al., 2024) or apply post-processing techniques to convert unstructured outputs into the desired format (Chen and Yang, 2023; Roegiest et al., 2023; Li et al., 2024). By eliminating this step, structured output enhances user experience and facilitates seamless engineering integration of LLMs into industrial applications such as agents (Liu et al., 2024b; Tam et al., 2024; Ayala and Bechard, 2024; Yuan, 2025).
While structured output is widely adopted in industrial scenarios (Ayala and Bechard, 2024), its impact on the generation quality of LLMs, compared to the standard unstructured format, remains insufficiently explored. Figure 1 presents a real-world example from the Last Letter Concatenation task (Wei et al., 2022), showing the inconsistent accuracies of different structured formats on the same case. Several prior studies have examined this problem but tend to reach one-sided conclusions that structured output either improves or impairs generation quality. Guo et al. (2025) identified two common challenges in medical question answering: unstructured outputs often contain factual inaccuracies and incomplete information, which are not fully resolved by the Chain of Thought (CoT) strategy (Wei et al., 2022). To address this, they propose explicitly structuring the output generation process of LLMs, resulting in more factual and comprehensive responses. Tam et al. (2024) present a different perspective, reporting a decline in the reasoning performance of LLMs when generating structured outputs compared to free-form responses. However, Kurt challenges the conclusions in Tam et al. (2024) by conducting experiments on the same datasets and deriving that structured generation improves LLMs’ generation. This study attributes the contradictory findings to an experimental design in the earlier work: the failure to design comparable prompts for structured and unstructured formats, rendering the comparisons nonequivalent.
@@ 2 Methodology
Given the applications involving compliance and regulatory requirements in highly regulated industry, it is essential to conduct a detailed impact analysis. In our study, the adopted causal inference involves randomized controlled trials (RCTs), observational data collection, and statistical tests to determine the impact of structured output format on LLMs’ generation (Pearl and Mackenzie, 2018). In this section, we first present the variables involved in our analysis: instruction, output format, and LLMs’ generation. The first two serve as potential causal factors, while the latter represents the outcome of interest. We then discuss the interventions targeting the two causal factors. Next, we elaborate on the potential causal structures using directed acyclic graphs (DAGs), which we aim to validate across various reasoning tasks. Finally, we outline the overall causal discovery. Our method is inspired by Bao et al. (2025) that explores CoT’s impact on LLMs’ generation through causal inference, with substantial modifications for our context.
@@ 2.1 Involved variables
Instruction encompasses elements necessary to prompt LLMs’ generation, including persona definition (system prompt), contextual information, problem formulation, and illustrative in-context learning cases (Swanson et al., 2021).
@@ 2.2 Interventions of instruction
To manage experimental complexity, we follow Bao et al. (2025), focus on zero-shot learning, do not incorporate external contexts such as knowledge graphs or web search tools, and limit alternative instruction in the main text to paraphrasing the system prompt into diverse persona definitions (Appendix B). Additional interventions on context description and problem formulation, along with the results, are reported in Appendix G.
@@ 2.3 Interventions of output format
In addition to the straightforward strategy, which is compatible with any LLM for obtaining structured output, we also assess JSON output via function calling, as detailed in the Appendix E.
@@ 3.1 Datasets
We evaluate LLMs’ capabilities across different reasoning types using 7 datasets: Grade School Math 8K (GSM8K) (Cobbe et al., 2021) for mathematics, LLC for symbolic transformation (Wei et al., 2022), Shuffled Objects Tracking (SOT) for symbolic relation (Ghazal et al., 2013), German Credit Factuality (GCF) (Yuan et al., 2025) for factual verification, German Credit Causality (GCC) (Yuan et al., 2025) for financial inference, Operations-oriented Evaluation (OpsEval) for system optimization and fault analysis (Liu et al., 2025), and Execution-based Code Evaluation (XCodeEval) for code compilation (Khan et al., 2024). We also develop a new dataset, ELLC [^2] https://github.com/Han-Yuan-Med/structured-output/tree/main/ellc_dataset, which extends the original LLC task. ELLC requires LLMs to first perform symbolic transformation to select correct letters, followed by linguistic reasoning to infer the sequence of selected letters to form a valid word. Exemplificative cases, development steps, and benchmark results of ELLC are detailed in Appendix A and I.
@@ 3.2 Models
We evaluate GPT-4o and several small language models (SLMs), including GPT-oss-20B (GPT-oss, ), Llama-3.1-8B (Touvron et al., 2023), Phi-3.5-mini (Abdin et al., 2024), and Gemma-2-9B (Mesnard et al., 2024) in the impact analysis of structured output. For reproducibility, we release analytical notebooks with results [^3] https://github.com/Han-Yuan-Med/structured-output.
@@ 3.3 Results
GPT-4o demonstrates robust insensitivity to structured output formats. In 43 out of 48 causal structures (DAGs), structured outputs have no statistically significant effect on generation. However, this robustness is not universal. In the remaining 5 DAGs, 3 are CwoM, where both instruction and output format causally influence the model’s output. The experimental results highlight the necessity of the used multifaceted analysis that should not be simplified to a single-factor perspective or to coarse comparisons of aggregated metrics.
@@
SLMs are more sensitive to structured output formats. As shown in Appendix F, 7 out of 24 DAGs exhibit causal effects, compared to 5 out of 48 for GPT-4o. Also, different from GPT-4o’s 100% success rate across all datasets, the success rate of SLMs can drop below 50% in certain reasoning tasks. The differences in causal discovery results between GPT-4o and SLMs stem from a combination of structured output generation and reasoning-based problem-solving capabilities, both of which are strongly influenced by model scale and architectural design. As evidenced by GPT-oss-20B, the largest model among all SLMs, it demonstrates greater robustness to interventions in both instruction and output format.
Derived DAGs are robust under additional interventions and API calls. Specifically, additional interventions target context and problem formulation, yielding the same DAGs as those presented in Table 1 (Appendix G). However, when conditioning on formats, we find that GPT-4o is more robust to changes in persona than to changes in context or formulation. Furthermore, despite GPT-4o API non-determinism, multiple trials yield consistent DAGs (Appendix H). For SLMs, consistency is ensured by configuring the decoding mechanism to greedy search.
@@ 4 Discussion and Conclusion
First, we acknowledge that some LLMs, such as Qwen (Yang et al., 2024) and DeepSeek (Liu et al., 2024a). Although we are unable to include these experiments, our method is broadly generalizable and is not constrained to a specific foundation model or module. Second, additional techniques, such as few-shot learning or reflection (Wang et al., 2024; Zhang et al., 2024; Gandhi et al., 2024; Guan et al., 2025; Chen et al., 2025; Qi et al., 2025), can enhance the performance of LLMs. Considering the complexity introduced by additional influential factors and the uncertainty caused by selected in-context cases or reflection mechanisms, we keep this direction as part of our future work. Third, we use answer accuracy as a proxy for generation quality. However, Wang et al. (2024) have shown that LLMs can occasionally arrive at correct answers despite flawed reasoning. Future work will therefore examine the impact of structured output on LLMs’ reasoning processes. Lastly, structured output can be achieved through other methods, such as output parsers, format-specific modes, etc. Our study focuses on format-restricting instructions and function calling because of their broad compatibility with most LLMs in industrial practice.
@@ Appendix A Datasets and LLC development
Building on LLC, we introduce a more complex task called ELLC. While LLC requires static concatenation of the last letters of given words, ELLC adds a linguistic reasoning component: models must determine the correct sequence of these letters to form a valid English word. We construct two versions: one using four input words, and another using six input words. Each sample includes a list of valid target words, as the same set of extracted letters can correspond to one or multiple correct answers. A response is considered correct if it matches any word in the target list. More challenging versions of this task involve a larger set of candidate words and letters, more complex extraction patterns, and the requirement for the model to generate all valid English words from the extracted letters. The main text reports experimental results for the basic task, while results for more complex variants are provided in the Appendix I.
@@
GCF and GCC are two datasets designed to assess LLMs’ ability to critically reflect on content generated by LLMs. This task is not easy, even for GPT-4o. A potential reason is that LLMs, despite being developed by different research institutions or companies, are often trained on overlapping corpora, such as Wikipedia. This overlap can result in high conditional probabilities being assigned to flawed outputs, making them appear more credible to LLMs than they do to humans (Gao et al., 2023). Both datasets consist of the 1,782 reasoning points produced by LLMs in the context of financial decision-making, each annotated for the presence of factual or causal issues (Wu et al., 2025). All samples are included in our experiments.
@@ Appendix E Function calling
Appendix F Small language models

### v2

# Quantifying the Impact of Structured Output Format on Large Language Models through Causal Inference

@@ Abstract
Structured output from large language models (LLMs) has enhanced efficiency in processing generated information and is increasingly adopted in industrial applications. Prior studies have investigated the impact of structured output on LLMs’ generation quality, often presenting one-way findings. Some suggest that structured format enhances completeness and factual accuracy, while others argue that it restricts the reasoning capacity of LLMs and leads to reductions in standard evaluation metrics. Potential limitations of these assessments include restricted testing scenarios, weakly controlled comparative settings, and reliance on coarse metrics. In this work, we present a refined analysis using causal inference. Based on one assumed and two guaranteed constraints, we derive five potential causal structures characterizing the influence of structured output on LLMs’ generation: (1) collider without m-bias, (2) collider with m-bias, (3) single cause from instruction, (4) single cause from output format, and (5) independence. Across seven public and one developed reasoning tasks, we find that coarse metrics report positive, negative, or neutral effects of structured output on GPT-4o’s generation. However, causal inference reveals no causal impact in 43 out of 48 scenarios. In the remaining 5, 3 involve multifaceted causal structures influenced by concrete instructions. Further experiments show that OpenAI-o3 are more resilient to output formats than general-purpose GPT-4o and GPT-4.1, highlighting an unaware advantage of reasoning models.
Quantifying the Impact of Structured Output Format on Large Language Models through Causal Inference
@@ 1 Introduction
Structured output, defined as organizing generated content into a specific schema (Guo et al., 2025; OpenAI, ), has been increasingly supported by LLMs, enabling them to directly output information in user-specified formats such as JSON, XML, and YAML [^1] JSON: JavaScript Object Notation; XML: Extensible Markup Language; YAML: Yet Another Markup Language (Tang et al., 2024; Wu et al., 2024; Guo et al., 2024). This development addresses a key limitation of earlier LLMs, which required users to either fine-tune models for structured generation (Chan et al., 2024; Tang et al., 2024) or apply post-processing techniques to convert unstructured outputs into the desired format (Chen and Yang, 2023; Roegiest et al., 2023; Li et al., 2024b). By eliminating this step, structured output enhances user experience and facilitates seamless engineering integration of LLMs into industrial applications such as Model Context Protocol (MCP) for agentic LLMs (Liu et al., 2024b; Tam et al., 2024; Ayala and Bechard, 2024; Yuan, 2025).
While structured output is widely adopted in industrial scenarios (Ayala and Bechard, 2024), its impact on the generation quality of LLMs, compared to the standard unstructured format, remains insufficiently explored. Figure 1 presents a real-world example from the Last Letter Concatenation task (Wei et al., 2022), showing the inconsistent accuracies of different structured formats on the same case. Several prior studies have examined this problem but tend to reach one-sided conclusions that structured output either improves or impairs generation quality. Guo et al. (2025) identified two common challenges in medical question answering: unstructured outputs often contain factual inaccuracies and incomplete information, which are not fully resolved by the Chain of Thought (CoT) strategy (Wei et al., 2022). To address this, they propose explicitly structuring the output generation process of LLMs, resulting in more factual and comprehensive responses. Tam et al. (2024) present a different perspective, reporting a decline in the reasoning performance of LLMs when generating structured outputs compared to free-form responses. Similar observations regarding the negative impact of structured output on generation quality have been reported in recent studies (Wang et al., 2025; Jang and Silavong, 2025; Chen et al., 2025a; Shi et al., 2025). However, Kurt challenges the conclusions in Tam et al. (2024) by conducting experiments on the same datasets and deriving that structured generation improves LLMs’ generation. This study attributes the contradictory findings to an experimental design in the earlier work: the failure to design comparable prompts for structured and unstructured formats, rendering the comparisons nonequivalent.
@@ 2 Methodology
Given the applications involving compliance and regulatory requirements in highly regulated industry, it is essential to conduct a detailed impact analysis. In our study, the adopted causal inference involves randomized controlled trials (RCTs), observational data collection, and statistical tests to determine the impact of structured output format on LLMs’ generation (Pearl and Mackenzie, 2018). In this section, we first present the variables involved in our analysis: instruction, output format, and LLMs’ generation. The first two serve as potential causal factors, while the latter represents the outcome of interest. We then discuss the interventions targeting the two causal factors. Next, we elaborate on the potential causal structures using directed acyclic graphs (DAGs), which we aim to validate across various reasoning tasks. Finally, we outline the overall causal discovery. Our method is inspired by Bao et al. (2025) that explores CoT’s impact on LLMs’ generation through causal inference, with substantial modifications for our context. Practitioners can refer to our released analytical notebook  [^2] https://github.com/Han-Yuan-Med/structured-output/blob/main/sot/sot_json.ipynb for detailed annotations and implementation procedures of the causal inference.
@@ 2.1 Involved variables
Instruction encompasses elements necessary to prompt LLMs’ generation, including persona definition (system prompt), contextual information, problem formulation, and in-context learning cases (Swanson et al., 2021; Miehling et al., 2025).
@@ 2.2 Interventions of instruction
To manage experimental complexity, we follow Bao et al. (2025), focus on zero-shot learning, do not incorporate external contexts such as knowledge graphs or web search tools, and limit alternative instruction in the main text to paraphrasing the system prompt into diverse persona definitions (Appendix B), which is a practical and used method of instruction perturbation in latest studies (Hu and Collier, 2024; Dong et al., 2024; Wu et al., 2025a; Li et al., 2025). Additional interventions on context description and problem formulation, along with the results, are reported in Appendix H.
@@ 2.3 Interventions of output format
In addition to the straightforward strategy, which is compatible with any LLM for obtaining structured output, we also assess JSON output via function calling, as detailed in the Appendix E. It is worth noting that, except for the experiments in Appendix E, all other experiments employ the format-restricting instruction strategy. Code for both format-restricting instruction and function calling settings is available on GitHub [^3] https://github.com/Han-Yuan-Med/structured-output/tree/main/generation.
@@ 3.1 Datasets
We evaluate LLMs’ capabilities across different reasoning types using 7 datasets: Grade School Math 8K (GSM8K) (Cobbe et al., 2021) for mathematics, LLC for symbolic transformation (Wei et al., 2022), Shuffled Objects Tracking (SOT) for symbolic relation (Ghazal et al., 2013), German Credit Factuality (GCF) (Yuan et al., 2025) for factual verification, German Credit Causality (GCC) (Yuan et al., 2025) for financial inference, Operations-oriented Evaluation (OpsEval) for system optimization and fault analysis (Liu et al., 2025), and Execution-based Code Evaluation (XCodeEval) for code compilation (Khan et al., 2024). We also develop a new dataset, ELLC [^4] https://github.com/Han-Yuan-Med/structured-output/tree/main/ellc_dataset, which extends the original LLC task. ELLC requires LLMs to first perform symbolic transformation to select correct letters, followed by linguistic reasoning to infer the sequence of selected letters to form a valid word. Exemplificative cases, development steps, and benchmark results of ELLC are detailed in Appendix A and J.
@@ 3.2 Models
We evaluate GPT-4o, GPT-4.1, OpenAI-o3, and several small language models (SLMs), including GPT-oss-20B (GPT-oss, ), Llama-3.1-8B (Touvron et al., 2023), Phi-3.5-mini (Abdin et al., 2024), and Gemma-2-9B (Mesnard et al., 2024) in the impact analysis of structured output. The temperature is fixed at 0 to eliminate randomness, and all other parameters remain at their default values. Unless otherwise specified, experiments use GPT-4o as the backbone. Experiments using GPT-4.1, OpenAI o3, and SLMs are reported in Appendices F and G. For reproducibility, we release the model outputs, analytical notebooks, and resulting quantitative outcomes on GitHub [^5] https://github.com/Han-Yuan-Med/structured-output.
@@ 3.3 Results
GPT-4o demonstrates robust insensitivity to structured output formats. In 43 out of 48 causal structures (DAGs), structured outputs have no statistically significant effect on generation. However, this robustness is not universal. In the remaining 5 DAGs, 3 are CwoM, where both instruction and output format causally influence the model’s output. The experimental results highlight the necessity of the used multifaceted analysis that should not be simplified to a single-factor perspective or to coarse comparisons of aggregated metrics. GPT-4.1 follows trends similar to GPT-4o, performing robustly in most settings but still exhibiting vulnerability on the SOT task. By contrast, OpenAI-o3 is consistently resilient to structured outputs, showing an underappreciated strength of reasoning-intensive models to general-purpose models.
@@
SLMs are more sensitive to structured output formats. As shown in Appendix G, 7 out of 24 DAGs exhibit causal effects, compared to 5 out of 48 for GPT-4o. Also, different from GPT-4o’s 100% success rate across all datasets, the success rate of SLMs can drop below 50% in certain reasoning tasks. The differences in causal discovery results between GPT-4o and SLMs stem from a combination of structured output generation and reasoning-based problem-solving capabilities, both of which are strongly influenced by model scale and architectural design. As evidenced by GPT-oss-20B, the largest model among all SLMs, it demonstrates greater robustness to interventions in both instruction and output format.
Derived DAGs are robust under additional interventions and API calls. Specifically, additional interventions target context and problem formulation, yielding the same DAGs as those presented in Table 1 (Appendix H). However, when conditioning on formats, we find that GPT-4o is more robust to changes in persona than to changes in context or formulation. Furthermore, despite GPT-4o API non-determinism, multiple trials yield consistent DAGs (Appendix I). For SLMs, consistency is ensured by configuring the decoding mechanism to greedy search.
@@ 4 Discussion and Conclusion
First, we acknowledge that some LLMs, such as Qwen (Yang et al., 2024) and DeepSeek (Liu et al., 2024a). Although we are unable to include these experiments, the introduced method is not constrained to a specific foundation model or module and is straightforward to apply to additional models seamlessly, offering a practical diagnostic tool for practitioners quantifying the impact of structured output formats. Second, additional techniques, such as few-shot learning or reflection (Wang et al., 2024; Zhang et al., 2024; Gandhi et al., 2024; Guan et al., 2025; Chen et al., 2025b; Qi et al., 2025), can enhance the performance of LLMs. Considering the complexity introduced by additional influential factors and the uncertainty caused by selected in-context cases or reflection mechanisms, we keep this direction as part of our future work. Third, we use answer accuracy as a proxy for generation quality. However, Wang et al. (2024) have shown that LLMs can occasionally arrive at correct answers despite flawed reasoning. Future work will therefore examine the impact of structured output on LLMs’ reasoning processes. Lastly, structured output can be achieved through other methods, such as output parsers, format-specific modes, etc. Our study focuses on format-restricting instructions and function calling because of their broad compatibility with most LLMs in industrial practice.
@@ Appendix A Datasets and LLC development
Building on LLC, we introduce a more complex task called ELLC. While LLC requires static concatenation of the last letters of given words, ELLC adds a linguistic reasoning component: models must determine the correct sequence of these letters to form a valid English word. We construct two versions: one using four input words, and another using six input words. Each sample includes a list of valid target words, as the same set of extracted letters can correspond to one or multiple correct answers. A response is considered correct if it matches any word in the target list. More challenging versions of this task involve a larger set of candidate words and letters, more complex extraction patterns, and the requirement for the model to generate all valid English words from the extracted letters. The main text reports experimental results for the basic task, while results for more complex variants are provided in the Appendix J.
@@
GCF and GCC are two datasets designed to assess LLMs’ ability to critically reflect on content generated by LLMs. This task is not easy, even for GPT-4o. A potential reason is that LLMs, despite being developed by different research institutions or companies, are often trained on overlapping corpora, such as Wikipedia. This overlap can result in high conditional probabilities being assigned to flawed outputs, making them appear more credible to LLMs than they do to humans (Gao et al., 2023). Both datasets consist of the 1,782 reasoning points produced by LLMs in the context of financial decision-making, each annotated for the presence of factual or causal issues (Wu et al., 2025b). All samples are included in our experiments.
@@ Appendix D Bonferroni correction
Hypotheses Fixed modules Examined modules SOT OpsEval XCodeEval
A
Standard & alternative
instruction
Controlled
(w/ standard format)
0.985 0.280 0.845
Intervened
(w/ JSON format)
0.859
0.265
0.842
Intervened
(w/ XML format)
0.961
0.265
0.855
Intervened
(w/ YAML format)
0.934
0.290
0.849
B Standard & JSON format
Controlled
(w/ standard instruction)
0.945 0.300 0.848
Intervened
(w/ alternative instruction)
0.916
0.266
0.843
Standard & XML format
Controlled
(w/ standard instruction)
0.985 0.275 0.850
Intervened
(w/ alternative instruction)
0.970
0.272
0.850
Standard & YAML format
Controlled
(w/ standard instruction)
0.975 0.300 0.848
Intervened
(w/ alternative instruction)
0.956
0.281
0.847
C LLMs’ generation JSON format (0.256) - -
XML format (0.710) - -
YAML format (0.435) - -
Derived DAGs JSON format CwoM, CwoM IND, IND IND, IND
XML format INS, CwoM IND, IND IND, IND
YAML format CwoM, CwoM IND, IND IND, IND
Table 4: Discovery of GPT-4.1’s DAGs based on structured output by format-restricting instruction. The $p$-values are reported in parentheses, with 0.000 indicating a value smaller than 0.0005 due to rounding to three decimal places. Each setting shows two DAGs based on $\alpha$ values of 0.05 (first) and 0.1 (second).
Hypotheses Fixed modules Examined modules SOT OpsEval XCodeEval
A
Standard & alternative
instruction
Controlled
(w/ standard format)
0.988 0.305 0.854
Intervened
(w/ JSON format)
1.000
0.310
0.858
Intervened
(w/ XML format)
1.000
0.295
0.847
Intervened
(w/ YAML format)
1.000
0.315
0.862
B Standard & JSON format
Controlled
(w/ standard instruction)
1.000 0.325 0.860
Intervened
(w/ alternative instruction)
0.999
0.303
0.855
Standard & XML format
Controlled
(w/ standard instruction)
1.000 0.312 0.860
Intervened
(w/ alternative instruction)
0.999
0.297
0.848
Standard & YAML format
Controlled
(w/ standard instruction)
1.000 0.325 0.865
Intervened
(w/ alternative instruction)
0.999
0.306
0.856
C LLMs’ generation JSON format - - -
XML format - - -
YAML format - - -
Derived DAGs JSON format IND, IND IND, IND IND, IND
XML format IND, IND IND, IND IND, IND
YAML format IND, IND IND, IND IND, IND
Table 5: Discovery of OpenAI-o3’s DAGs based on structured output by format-restricting instruction. The $p$-values are reported in parentheses, with 0.000 indicating a value smaller than 0.0005 due to rounding to three decimal places. Each setting shows two DAGs based on $\alpha$ values of 0.05 (first) and 0.1 (second).
@@ Appendix E Function calling
Appendix F General-purpose and reasoning models
Most experiments in the main text use the general-purpose GPT-4o as the backbone. To further examine the impact of structured output on the latest general-purpose and reasoning models, we additionally conduct experiments on GPT-4.1 and OpenAI-o3. Tables 4 and 5 presents the DAGs of these two models released after GPT-4o. GPT-4.1, the general-purpose successor, exhibits patterns similar to GPT-4o, indicating that its performance on the SOT dataset is statistically significantly affected by structured output formats. In contrast, OpenAI-o3, a high-performance reasoning model, achieves perfect accuracy on the SOT dataset across all settings except for a single case involving intervened instruction under the unstructured format. Consequently, its DAGs and p-values remain stable across structured formats. On the other two datasets of OpsEval and XCodeEval, the content quality of OpenAI-o3 is also not statistically affected by output format, highlighting a complementary advantage of reasoning-intensive models over general-purpose models.
Appendix G Small language models